# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ III.2 Oscillatory Drive
In Fig. 5, we revisit the stroboscopic time evolution for $\widehat{N}(t)$ at the numerically determined renormalized freezing points (Fig. 4). For |$\ket{\psi(0)}$, we choose both the ground state and an excited eigenstate of $H(0)$; $\widehat{N}$ becomes approximately conserved at the renormalized freezing points. The slow time decay associated with the emergent conservation in Fig. 5(c) is controlled by the higher order corrections in the Magnus expansion.
@@
Finally, in Fig. 6 we show the stroboscopic evolution of the half-chain entanglement entropy. At the renormalized freezing points (indicated by red curves), there is a noticeable slow increase in the entanglement entropy compared to the square-wave drive. The difference stems primarily from the structure of higher-order corrections in a Magnus expansion; while the first-order correction for the square-wave drive is zero, it is non-zero for the oscillatory drive. However, away from freezing (represented by the blue curves), the system rapidly thermalizes, and the difference from the frozen curve is striking.
Figure 4: (a) Time-averaged ${\cal{F}}$ (averaged over 100 Floquet cycles) as a function of $A/\omega$ for the cosine drive. We choose |$\ket{\psi(0)}$ from the eigenspectrum of $H(0)$, with their “state-index” sorted by energy eigenvalue. Primary freezing points, indicated by arrows, shift from the zeroth-order prediction of freezing point in Eq. 14b (dahsed lines). The shift can be explained based on the higher order corrections in the Magnus expansion, discussed in Sec. V. (b) The eigenspectrum of the Hamiltonian $H(0)$ is plotted along with their corresponding Fock state indices. The states chosen to calculate $\langle\widehat{N}(t)\rangle$ are indicated with arrows.
Figure 5: Stroboscopic time evolution of normalized $\widehat{N}(t)$ for the cosine drive given by Eq.(3b). The system size is $L=8$, with an on-site Hilbert space dimension $\mathcal{H}_{D}=3$. The drive frequency $\omega=9.32$ GHz, and other parameters are as introduced in the main text. The initial state are chosen from the eigenstates of $H(0)$. The renormalized freezing points are obtained from Fig. 4.
Figure 6: Stroboscopic evolution of half-chain entanglement entropy at (red), and away from (blue) freezing for a cosine wave drive for the parameters in Fig. 5 for different initial states, |$\ket{\psi(0)}$.
IV Results in the Semi-Classical Limit
@@
The coupled semi-classical equations of motion in Eq. 4a-4b for $\varphi_{i}$ and $n_{i}$ can be solved using a Tsitouras 5/4 Runge-Kutta method [68], starting from a specific initial state. One of our main goals in this section will be to analyze the chaotic properties associated with the resulting trajectories. Consider two trajectories in phase space with an infinitesimal initial separation $\delta\pi_{0}$ at time $t=0$, and let $\delta\pi(t)$ be the resulting separation at a later time $t$. The maximal Lyapunov exponent is defined as [69]
@@
In order to pick a starting state that has a close correspondence with realistic “computational states” in the quantum problem, we adapt the strategy highlighted in Ref. [46]. First we calculate the eigenenergies, $E_{m}$, for a single static quantum transmon. The energies of the computational states are given by $E_{0},~{}E_{1}$ for the $|0\rangle,~{}|1\rangle$ state, respectively. We then initialize the transmons in the semiclassical limit as 00 or $1$ by initializing their phase space coordinates as $(n,\varphi)$=$(0,\varphi_{m})$, where $-E_{J}\cos{\varphi_{m}}=E_{m}$ for $m=0,~{}1$. In this fashion we can initialize semi-classical states as $101010\cdots$, that mimic the corresponding quantum state. The Lyapunov spectrum is computed using the “H2 method”  [70, 71], and implemented in Julia  [72].
Figure 7: (a) The maximum Lyapunov exponent (MLE), $\lambda_{L}$, as a function of $A/\omega$ for a fixed drive frequency ($\omega=15$ GHz) with the square-wave drive, initialized in the $1010$ state. The strong suppression of $\lambda_{L}$ coincides with the location of freezing points in Eq. 14a. Dashed horizontal line shows $\lambda_{L}$ for the static hamiltonian. (b) Correspondence between quantum and classical dynamics as a function of the ratio of drive frequency to transmon frequency at freezing with $A/\omega=2$ kept fixed. The quantum expectation $\langle\widehat{N}(t)\rangle$ is computed by time-averaging over 100 Floquet cycles in the ground state of $H(0)$. For both plots $L=8$, $E_{c}=12.58$ GHz, $E_{J}=330$ MHz, and $V=50$ MHz.
@@
We begin first by discussing the results for the square-wave drive in Eq. 3a. In Fig. 7(a) we plot the MLE as a function of the drive amplitude, $A$, for a fixed drive frequency, $\omega$. Remarkably, whenever $A/\omega$ approaches the special points in Eq. 14a, $\lambda_{L}$ is suppressed by several orders in magnitude compared to the value in the static case (shown by dashed line). Relatedly, to highlight the correspondence between the quantum and classical results, we evaluate $\langle\widehat{N}(t)\rangle$ (normalized by its initial value) from the ED data, and $\lambda_{L}$ using the semiclassical equations as a function of the ratio of the drive to transmon frequencies, respectively. The results are shown in Fig. 7(b) at a fixed value of $A/\omega=2$, which corresponds to one of the freezing points. Clearly, both sets of results suggest that freezing requires a lower cutoff on the drive frequency, set roughly by the individual transmon frequency. We have also noticed that for a given value of $\omega$, freezing requires the amplitude to be larger than a critical value.
@@
To help visualize how the trajectories for $\{\varphi,~{}n\}$ evolve with varying $A/\omega$ for the square-wave drive, we generate phase-space portraits for two transmons (picked at random) from the total system consisting of $L=8$ transmons. When the system is away from freezing, and starting from an initial state $1010$, the phase-space fills up uniformly over time and displays ergodic behavior as shown in Fig. 8(a). Specifically, $n$ displays strong fluctuations over time starting from its initial value. In contrast, at freezing the individual transmons exhibit strong deviations away from any ergodic behavior (Fig. 8(b)), leading to a highly constrained set of trajectories in phase-space showing minimal fluctuations of $n$.
Figure 8: Phase-space trajectories for two transmons (site $2$ and $5$, respectively) for an array with $L=8$ transmons, for initial state $1010$ and a square-wave drive with $\omega=15$ GHz over $250$ cycles. The values of the drive amplitude are (a) $A=110$ GHz, and (b) $A=120.36$ GHz (corresponding to $A/\omega=8$), respectively.
@@
Finally, turning over to the cosine drive, Fig. 9 shows a plot of $\lambda_{L}$ as a function of $A/\omega$ for an $L=8$ transmon array. Once again, we find that $\lambda_{L}$ drops significantly at freezing compared to the static Hamiltonian. However, the location of the freezing points is not determined by the zeroth-order Magnus expansion, as was noted in the ED analysis as well, highlighting the significant role played by the higher order corrections.
Figure 9: The maximum Lyapunov exponent as a function of $A/\omega$ for the cosine drive. The system with $L=8$ transmons is initialized in an $1010$ state with a drive frequency $\omega=15$ GHz. The freezing points are indicated by arrows. The horizontal dashed line indicates $\lambda_{L}$ for the static Hamiltonian. The vertical dotted lines indicate the freezing points associated with the zeroth-order term in Eq. 14b.
V Floquet-Magnus Expansion
@@
In this section, we extend the Floquet-Magnus expansion for the transmon Hamiltonian beyond the leading order discussed previously in Eq. 13-13b. In particular, we use the expansion to compute the renormalized freezing point. As will become clear later, it will be useful to consider the drive amplitude to vary across different sites. Thus we modify the drive term as,
@@
where the $\theta_{i}$ is expressed as an integral over the $f_{i}$’s as in Eq. 8, but is position-dependent now.
The first few terms in the Magnus expansion of the Floquet Hamiltonian in terms of the co-moving Hamiltonian Eq. 10 are given by
@@
The integrals can be evaluated analytically (see Appendix  A), and we only summarize the results below. The zeroth-order term is,
@@
$H^{2}_{A}$ $=$ $\frac{E_{J}^{2}}{8\omega^{2}}\sum_{kl}\big{[}B_{k}^{*}B_{l}^{*}e^{i{\widehat{\varphi}}_{k}+i{\widehat{\varphi}}_{l}}\Delta_{k}\Delta_{l}$ (29a)
@@
$H^{2}_{B}$ $=$ $\frac{E_{J}^{2}}{4\omega^{2}}\sum_{kl}\left[C_{k}^{*}F_{0,l}^{*}e^{i{\widehat{\varphi}}_{k}+i{\widehat{\varphi}}_{l}}\Delta_{k}\Delta_{l}\right.$ (29b)
@@
The parameter $C_{k}$ is
@@
Next, we consider the first-order renormalizations of the freezing point, which is especially relevant for the cosine drive in Eq. 3b. We write the Floquet Hamiltonian as $\mathcal{H}_{\text{eff}}=\mathcal{H}^{0}_{\widehat{n}}+U$, where
@@
Eqs.(33) and (34) is the main result of this section, which indicates the first-order correction of the freezing point in terms of the amplitude-frequency ratio $A_{k}/\omega$. Notably, the right hand side of Eq. (34) is proportional to $n_{i}$. Thus, the magnitude of renormalization a state experiences from the zeroth order freezing value increases with increasing $n_{i}$.
Figure 10: The time-averaged observable, ${\cal{F}}$ (Eq. 16), as a function of $\omega$ rescaled by the individual transmon frequency, $\nu$. We choose |$\ket{\psi(0)}$ as the ground state of $H(0)$ and fix $A/\omega=2.4048$ at the first freezing point within the zeroth-order Magnus expansion. The scaling is predominantly proportional to $1/\omega$ (Eq. 25).
Figure 11: ${\cal{F}}$ as in Fig. 10, but at a value of $A/\omega$ which is set at the renormalized freezing point (taking into account the zeroth and first-order corrections). The parameters are the same as in Fig.10.
We notice, however, that satisfying Eq. (33) does not completely eliminate the first-order corrections. At the first-order corrected freezing point (33), there is still a residual action of $U$ on |$\ket{n}$:
@@
This term arises exactly due to the non-commutativity between $\Delta_{k}\mathcal{H}^{0}_{\widehat{n}}=4E_{c}(2\hat{n}_{k}+1)$ and $e^{i{\widehat{\varphi}}_{k}}$, which is the manifestation of uncertainty principle in the context of dynamical-freezing. We also note that the first-order renormalization depends on the initial state |$\ket{n}$ and can be spatially inhomogeneous. In Figs. 10 and 11, we test our numerical results with the theoretical prediction of the freezing under cosine drive. We compared the decay rate of $\overline{\langle N(t)\rangle}$ for $A/\omega$ at the zeroth-order freezing point and the first-order freezing point and performed a scaling analysis in $\omega$. In both cases the decay rate primarily scales as $1/\omega$. We observe that the decay rate of $\overline{\langle N(t)\rangle}$ is reduced at the first-order freezing point, in the sense that the coefficient of $1/\omega$ becomes smaller (but not vanishing). These behaviors agree with the theoretical prediction that the decay rate scales as $1/\omega$.
Next, we discuss the second-order corrections and focus on the square-wave drive Eq. 3a. Since $B_{k}=0$ (see Eq. 28), the first-order correction vanishes. Among the second-order corrections, $H_{A}^{2}=0$ due to $B_{k}=0$, $H_{B}^{2}\sim 1/\omega^{4}$ due to $F_{0}\sim 1/\omega^{2}$ and we therefore only need to consider $H^{2}_{C}$. We note that because $H^{2}_{C}$, similar to $H^{0}_{\text{eff}}$, depends on ${\widehat{\varphi}}$ only through single-site operators, it can be partially cancelled by adjusting $A_{k}/\omega$. The computation of the renormalized freezing point proceeds similarly to the first-order case, with the result
@@
Again, this does not completely cancel the second order correction due to non-commutativity, and we therefore expect the residual non-conservation rate to scales as $1/\omega^{2}$ for the system under square-wave drive.
We have performed ED computations including disordered couplings, $E_{J}$, drawn from a random distribution with a finite variance. For a small variance, the phenomenon of freezing persists. However, given our observation of the site-dependent corrections in the higher-order Magnus expansion, an increasing strength of disorder can affect the location of freezing. In the asymptotic sense, for any given strength of disorder, a strong enough drive frequency can be used to reinstate the freezing phenomenology in principle. However, we leave a detailed study of this for future work.
VI Outlook

### v2

@@ Results
In Fig. LABEL:fig:cosinendt(a),(b),(c), we revisit the stroboscopic time evolution for $\widehat{N}(t)$ at the numerically determined renormalized freezing points (Fig. LABEL:fig:cosonefidelity(a)). For |$\ket{\psi(0)}$, we choose both the ground state and an excited eigenstate of $H(0)$ (as shown by the colored arrows in Fig. LABEL:fig:cosonefidelity(b)); $\widehat{N}$ becomes approximately conserved at the renormalized freezing points. The slow time decay associated with the emergent conservation in Fig. LABEL:fig:cosinendt(c) is controlled by the higher order corrections in the Magnus expansion.
@@
Finally, in Fig. LABEL:fig:cosinendt(d),(e),(f) we show the stroboscopic evolution of the half-chain entanglement entropy. At the renormalized freezing points (indicated by red curves), there is a noticeable slow increase in the entanglement entropy compared to the square-wave drive. The difference stems primarily from the structure of higher-order corrections in a Magnus expansion; while the first-order correction for the square-wave drive is zero, it is non-zero for the oscillatory drive. However, away from freezing (represented by the blue curves), the system rapidly thermalizes, and the difference from the frozen curve is striking. The noticeable growth of entanglement entropy in this case even at the freezing point highlights the influence of interactions as was pointed out in Ref. 29.
The exact diagonalization computations have been performed here for a finite system and Hilbert-space dimension, respectively, both of which introduce an exponential complexity. We have studied the systematic effects of both of these finite-size corrections on the freezing phenomenology in Supplemental MethodsA. Additionally, incorporating a time evolution axis makes the numerics challenging. For larger systems, where exact diagonalization becomes unfeasible, alternative methods like tensor network techniques, such as the infinite time-evolving block decimation (iTEBD) algorithm, remain an exciting avenue for future studies.
Semi-Classical Limit
@@
The transmon system we consider is amenable to a semi-classical treatment where we take $\hat{\phi}$ and $\hat{n}$ as real variables rather than quantum operators; see methods section LABEL:sec:SClimit for details. In this section, we present results of this semi-classical analysis. The coupled semi-classical equations of motion in Eq. (41a)-(41b) for $\varphi_{i}$ and $n_{i}$ can be solved using a Tsitouras 5/4 Runge-Kutta method 65, starting from a specific initial state. Our main goals in this section will be to analyze the chaotic properties associated with the resulting trajectories and observe their dependence on system size and varying amounts of disorder.
@@
In order to pick a starting state that has a close correspondence with realistic “computational states” in the quantum problem, we adapt the strategy highlighted in Ref. 45. First we calculate the eigenenergies, $E_{m}$, for a single static quantum transmon. The energies of the computational states are given by $E_{0},~{}E_{1}$ for the $|0\rangle,~{}|1\rangle$ state, respectively. We then initialize the transmons in the semiclassical limit as 00 or $1$ by initializing their phase space coordinates as $(n,\varphi)$=$(0,\varphi_{m})$, where $-E_{J}\cos{\varphi_{m}}=E_{m}$ for $m=0,~{}1$. We have performed additional simulations that suggest that the semi-classical freezing phenomenology discussed below is only strengthened by initializing to non-zero values of $n$. In this fashion we study the evolution with an initial semi-classical state $101010\cdots$. The maximum Lyapunov exponent is computed using the “H2 method”  66, 67 and implemented in Julia  68.
Phenomenology of Semi-Classical Freezing
@@
We begin first by discussing the results for the square-wave drive in Eq. (3a). In Fig. LABEL:fig:classicalLE(a) we plot the MLE as a function of the drive amplitude, $A$, for a fixed drive frequency, $\omega$. Remarkably, whenever $A/\omega$ approaches the special points in Eq. (40a), $\lambda_{L}$ is suppressed by over an order of magnitude compared to the value in the static case (shown by dashed line). To further highlight the correspondence between the quantum and classical results, we analyze freezing as a function of the drive frequency, $\omega$, in both the quantum and semi-classical cases. The Magnus expansion, which can account for many of our numerical results, is formally a large $\omega$ expansion, so we might expect that there exists a threshold for $\omega$ below which freezing no longer occurs. To see this, we evaluate $\langle\widehat{N}(t)\rangle$ (normalized by its initial value) from the ED data, and $\lambda_{L}$ calculated from the semi-classical equations for the same system size and parameters. We plot both as a function of $\omega/\nu$ where $\nu$ is the fundamental transmon frequencies defined in Eq. (4). The results are shown in Fig. LABEL:fig:classicalLE(b) at a fixed value of $A/\omega=2$, which corresponds to one of the freezing points. Both sets of results suggest that freezing requires the drive frequency to be above a threshold set roughly by the individual transmon frequency (single particle energy scale). We have also noticed that for a given value of $\omega$, freezing requires the drive amplitude to also be larger than a critical value. For a more in depth discussion of the $\omega$ dependence of semi-classical freezing, see Supplemental Methods D.
@@
To help visualize how the trajectories for $\{\varphi,~{}n\}$ evolve with varying $A/\omega$ for the square-wave drive, we generate phase-space portraits for two transmons (picked at random) from the total system consisting of $L=8$ transmons. When the system is away from freezing, and starting from an initial state $1010$, the phase-space fills up uniformly over time and displays ergodic behavior as shown in Fig. LABEL:fig:phasespace8(a). Specifically, $n$ displays strong fluctuations over time starting from its initial value. In contrast, at freezing the individual transmons exhibit strong deviations away from any ergodic behavior (Fig. LABEL:fig:phasespace8(b)), leading to a highly constrained set of trajectories in phase-space showing minimal fluctuations of $n$.
@@
Finally, turning to the cosine drive, Fig. LABEL:Lyapnuov_cosinedrive shows a plot of $\lambda_{L}$ as a function of $A/\omega$ for an $L=8$ transmon array. Once again, we find that $\lambda_{L}$ drops significantly at freezing compared to the static Hamiltonian. However, the location of the freezing points is somewhat shifted from the predictions of the zeroth-order Magnus expansion. This was noted in the ED analysis as well, highlighting the significant role played by the higher order corrections in the case of the cosine drive.
Dependence on System Size and Disorder
In this section, we analyze the extent of dynamical freezing within our semi-classical setup in systems of up to $433$ transmons arranged in the same geometries that IBM uses for its existing quantum devices. Clearly, this lies beyond the range of exact diagonalization, but given the excellent correspondence between the two sets of computations for small system-sizes, demonstrates the likely fate of freezing for larger systems. We then turn to the effect of disorder in the Josephson energy coupling on freezing.
The particular two-dimensional “heavy-hexagon” geometries that IBM 62 uses in its currently available quantum devices are shown in Fig. LABEL:fig:IBM_simulations(a). For each of the Falcon, Hummingbird, Eagle, and Osprey devices, we initialize the semi-classical system in a checkerboard pattern of $1$’s and 00’s. In Fig. LABEL:fig:IBM_simulations(b), we plot the MLE as a function of the drive amplitude $A$ at a fixed frequency $\omega$. The freezing phenomenology discussed above persists for all devices considered with $\lambda_{L}$ being significantly suppressed at freezing points. Furthermore, we find that with increasing system size, the growth of $\lambda_{L}$ is larger away from the freezing points.
So far we have worked exclusively with translationally invariant Hamiltonians. However, realistic transmon arrays have disordered values of $E_{J_{i}}$, which are either intentionally engineered or due to unintended imperfections. To study the effect of disorder on dynamical freezing in the semi-classical limit, we consider a one-dimensional array of length $8$ with $E_{C}=330$ MHz, $V=50$ MHz, and $E_{J_{i}}\sim\mathcal{N}(12.58,\sigma^{2})$ GHz. Here, the values of $E_{J_{i}}$ are chosen from a normal distribution with standard deviation $\sigma$ 69, independently at each site. To ensure we stay in the transmon regime, we impose a cutoff of $4$ GHz as the minimum value of $E_{J_{i}}$. After driving the system with the square-wave drive at $\omega/(2\pi)=15$ GHz and calculating $\lambda_{L}$, we average over $1000$ disorder realizations. In Fig. LABEL:fig:Disorder_averaging we plot $\lambda_{L}$ as a function of $\sigma$ for $A/\omega=0,2.17,4.04$ – respectively the undriven case and the numerically-calculated renormalized values for the first and second freezing points. We find that for small values of $\sigma$, the phenomenon of freezing persists. However, much larger values of $\sigma$ weaken freezing (particularly for the second freezing point) while simultaneously reducing chaos in the undriven case (the latter behavior being in alignment with previous results 45). Nonetheless, we expect that in the asymptotic sense, for any given strength of disorder, a strong enough drive frequency can be used to reinstate the freezing phenomenology in principle. However, we leave a detailed study of this for future work.
Floquet-Magnus Expansion
@@
In this section, we extend the Floquet-Magnus expansion for the transmon Hamiltonian beyond the leading order, as described in the methods section LABEL:sec:FM. In particular, we use the expansion to compute the renormalized freezing point. As will become clear later, it will be useful to consider the drive amplitude to vary across different sites. Thus we modify the drive term as,
@@
where the $\theta_{i}$ is expressed as an integral over the $f_{i}$’s as in Eq. (34), but is position-dependent now.
The first few terms in the Magnus expansion of the Floquet Hamiltonian in terms of the co-moving Hamiltonian Eq. (36) are given by
@@
The integrals can be evaluated analytically (see Supplemental Methods summarize the results below. The zeroth-order term is,
@@
$H^{2}_{A}$ $=$ $\frac{E_{J}^{2}}{8\omega^{2}}\sum_{kl}\big{[}D_{kl}^{*}e^{i{\widehat{\varphi}}_{k}+i{\widehat{\varphi}}_{l}}\Delta_{k}\Delta_{l}\mathcal{H}^{0}_{\widehat{n}}$ (18a)
@@
$H^{2}_{B}$ $=$ $\frac{E_{J}^{2}}{4\omega^{2}}\sum_{kl}\left[C_{k}^{*}F_{0,l}^{*}e^{i{\widehat{\varphi}}_{k}+i{\widehat{\varphi}}_{l}}\Delta_{k}\Delta_{l}\mathcal{H}^{0}_{\widehat{n}}\right.$ (18b)
@@
The parameters $D_{kl}$, $E_{kl}$, and $C_{k}$ are
$D_{kl}=B_{k}B_{l}-\frac{1}{2}\sum_{m\neq 0}\frac{F_{-m,k}F_{m,l}}{m^{2}}\,,$ (19)
$E_{kl}=B_{l}B_{k}^{*}+\frac{1}{2}\sum_{m\neq 0}\frac{F_{m,k}^{*}F_{-m,l}}{m^{2}}\,,$ (20)
@@
Next, we consider the first-order renormalizations of the freezing point, which is especially relevant for the cosine drive in Eq. (3b). We write the Floquet Hamiltonian as $\mathcal{H}_{\text{eff}}=\mathcal{H}^{0}_{\widehat{n}}+U$, where
@@
Eqs. (24) and (25) are the main result of this section, which indicates the first-order correction of the freezing point in terms of the amplitude-frequency ratio $A_{k}/\omega$. Notably, the right hand side of Eq. (25) is proportional to $n_{i}$. Thus, the magnitude of renormalization a state experiences from the zeroth order freezing value increases with increasing $n_{i}$.
We notice, however, that satisfying Eq. (24) does not completely eliminate the first-order corrections. At the first-order corrected freezing point Eq. (24), there is still a residual action of $U$ on |$\ket{n}$:
@@
This term arises exactly due to the non-commutativity between $\Delta_{k}\mathcal{H}^{0}_{\widehat{n}}=4E_{c}(2\hat{n}_{k}+1)$ and $e^{i{\widehat{\varphi}}_{k}}$, which is the manifestation of uncertainty principle in the context of dynamical-freezing. We also note that the first-order renormalization depends on the initial state |$\ket{n}$ and can be spatially inhomogeneous. In Fig. LABEL:fig:zeroorderscaling we test our numerical results with the theoretical prediction of the freezing under cosine drive. We compared the decay rate of $\overline{\langle N(t)\rangle}$ for $A/\omega$ at the zeroth-order freezing point and the first-order freezing point and performed a scaling analysis in $\omega$. In both cases the decay rate primarily scales as $1/\omega$. We observe that the decay rate of $\overline{\langle N(t)\rangle}$ is reduced at the first-order freezing point, in the sense that the coefficient of $1/\omega$ becomes smaller (but not vanishing). These behaviors agree with the theoretical prediction that the decay rate scales as $1/\omega$.
Next, we discuss the second-order corrections and focus on the square-wave drive Eq. (3a). Since $B_{k}=0$ (see Eq. (17)), the first-order correction vanishes. Among the second-order corrections, $H_{B}^{2}\sim 1/\omega^{4}$ due to $F_{0}\sim 1/\omega^{2}$ and we therefore need to consider $H^{2}_{A}$ and $H^{2}_{C}$. We note that because $H^{2}_{C}$, similar to $H^{0}_{\text{eff}}$, depends on ${\widehat{\varphi}}$ only through single-site operators, it can be partially cancelled by adjusting $A_{k}/\omega$. The computation of the renormalized freezing point proceeds similarly to the first-order case, with the result
@@
Again, this does not completely cancel the second order correction due to non-commutativity.
Furthermore, at the freezing point $H^{2}_{A}\neq 0$, and it contains terms of the form $e^{i\hat{\varphi}_{k}\pm i\hat{\varphi}_{l}}$, which cannot be canceled by adjusting $A_{k}/\omega$. Therefore, both $H^{2}_{A}$ and $H^{2}_{C}$ will contribute to non-conservation at the freezing point, and we expect the residual non-conservation rate to scale as $1/\omega^{2}$ for the system under square-wave drive.
Discussion